OpenStudy (anonymous):

X/x^2-4 - 1/ x+ 2 = 2

OpenStudy (anonymous):

Solve please!

Parth (parthkohli):

\[\rm {x \over x^2 - 4} - {1 \over x + 2} = 2\]Factor the required stuff.\[\rm {x\over (x + 2)(x - 2)} - {1 \over x + 2} = 2\]Multiply -1/(x + 2)'s numerator and denominator by (x - 2)\[\rm {x \over (x + 2)(x - 2)}-{1 \over (x + 2)(x - 2)} = 2\]\[\rm {x - 1 \over (x + 2)(x - 2)} = 2\]

Parth (parthkohli):

To clear out the fractions, multiply both sides by \(\rm x - 1 = 2(x + 2)(x - 2) \)

OpenStudy (anonymous):

What do you mean by rm?